Hi I am a granfather of 75 yrs old. trying to help out his grandaughter. she has been working on a algebra problem for hours finish all but one question. Sent her home with her mother who tried also.she was so tired I sent her to bed. Its been 60 yrs since I had to do math..lol She just doesn't understand how to begin..Honestly we tried. if you can help it would be wonderful. the question is
two trains traveling in different directions one at 55mph and the other at 45mph
they traveled a total of 400 miles how long did it take to go that distance...
she doesn't know how to work out the formula...rate and time. thank you

Answer by stanbon(75887)

two trains traveling in different directions one at 55mph and the other at 45mph
they traveled a total of 400 miles how long did it take to go that distance...
-----------------------
The basic formula is: distance = rate * time
-------------------------
1st train DATA:
rate = 55 mph ; distance = x miles ; time = d/r = x/55 hrs.
---------------------
2nd train DATA:
rate = 45 mph ; distance = (400-x) miles ; time = d/r = (400-x)/45 hrs
------------------------
EQUATION:
time = time
x/55 = (400-x)/45
Cross-multiply to get:
45x = 55(400-x)
45x = 55*400 - 55x
100x = 55*400
x = 55*4
x = 220 miles
----------------
time = distance/rate
time = 220 / 55 = 4 hours
The time is the same for both of the trains.
==================
Another way to do the problem.
The trains are going away from one another at (45+55) = 100 mph
The time for them to be 400 miles apart is 400/100 = 4 hours.
=======================
Cheers,
Stan H.